Besty a recent retiree,requires $5,000 per year in extra income. She has $70,000 to invest and can invest in B-rated bonds paying 15% per year or in a certificate of deposit (CD) paying 7% a year. How much should be invested in each to realize exactly $5,000 in interest per year?
Answer by stanbon(75887)

Betsy a recent retiree,requires $5,000 per year in extra income. She has $70,000 to invest and can invest in B-rated bonds paying 15% per year or in a certificate of deposit (CD) paying 7% a year. How much should be invested in each to realize exactly $5,000 in interest per year?
-------------------------------------
Equation:
interest + interest = interest
---
0.15x + 0.07(70000-x) = 5000
Multiply thru by 100 to get:
15x + 7*70000-7x = 500000
-----
8x = 500000-7*70000
8x = 10000
x = $1,250.00 (amt. invested at 15%)
--------------
70000-x = $68750.00 (amt. invested at 7%)
=============================================
Cheers,
Stan H.